Compare three student loans for $80,000 for 4 years of college. Compare varying student loan offers, their monthly payments and total repayment cost.

Three possible examples:

Student 1) immediately. A loan at fixed 6.25% that you pay for 10 years starting right away as a normal installment loan,

Student 2) Pay interest only. Pay interest at fixed 6.25% for the four years you are in college. Then for 6 years after college, pay the loan as a 6 year installment loan.

Student 3) Defer your payment. Pay nothing while you are in college. After you graduate, pay the loan back as a 10 year loan at 8.25%. Interest rates are higher for this option.
Answer #1
1) PMT (using formula for loan amortization) = 80000*(0.0625/12)*(1+0.0625/12)^120/((1+0.0625/12)^120-1)) = $               898
Total amount payable = 898*120 = $      1,07,760
Total interest payable = 107760-80000 = $         27,760
2) Interest for 4 years = 80000*6.25%*4 = $         20,000
PMT for 6 years (using formula for loan amortization) = 80000*(0.0625/12)*(1+0.0625/12)^72/((1+0.0625/12)^72-1)) = $            1,335
Total installment payments = 1335*72 = $         96,120
Interest in the above = 96120-80000 = $         16,120
Total interest payable = 20000+16120 = $         36,120
Total amount payable = 20000+96120 = $      1,16,120
3) PMT (using formula for loan amortization) = 80000*(0.0825/12)*(1+0.0825/12)^120/((1+0.0825/12)^120-1)) = $               981
Total amount payable = 981*120 = $      1,17,720
Total interest payble = 117720-80000 = $         37,720
[It is assumed that no interest would be charged on the loan for
the 4 years of deferral; that is the loan outstanding at EOY 4
would be $80000]
4) The results are tabulated below:
Loan PMT Total Payment Total Interest
1 $               898 $ 1,07,760 $        27,760
2 $            1,335 $ 1,16,120 $        36,120
3 $               981 $ 1,17,720 $        37,720